Show the complete solution on the following questions.


1. Evaluate "\nabla \cdot (F \times G)," given "𝐹(𝑥,𝑦,𝑧) = 𝑦𝑧 \vec{𝑖} + 𝑥𝑧\vec{ 𝑗} + 𝑥𝑦\vec{ 𝑘}" and "𝐺(𝑥,𝑦,𝑧) = 𝑥𝑦 \vec{𝑗} + 𝑥𝑦𝑧\vec{ 𝑘}"


1
Expert's answer
2020-03-30T10:25:05-0400

"F×G=(F_yG_z-F_zG_y)\vec{i}+(F_zG_x-F_xG_z)\vec{j}+(F_xG_y-F_yG_x)\vec{k}"

"F×G=(x^2yz^2-x^2y^2)\vec{i}+(-xy^2z^2)\vec{j}+(xy^2z)\vec{k}"

"\nabla⋅(F×G)={\frac \partial {\partial x}}(x^2yz^2-x^2y^2)+{\frac \partial {\partial y}}(-xy^2z^2)+{\frac \partial {\partial z}}(xy^2z)"

"\nabla⋅(F×G)=2xyz^2-2xy^2-2xyz^2+xy^2=-xy^2"